The brakes of a car whose initial velocity is 30 m/s are applied, and the car receives an acceleration of −2 m/s2. How far will it have gone (a) when its velocity has decreased to 15 m/s and (b) when it has come to a stop?



1
Expert's answer
2021-11-01T12:53:30-0400

(a)

"v^2=v_0^2+2ad,""d=\dfrac{v^2-v_0^2}{2a},""d=\dfrac{(15\ \dfrac{m}{s})^2-(30\ \dfrac{m}{s})^2}{2\times(-2.0\ \dfrac{m}{s^2})}=169\ m."

(b)

"v^2=v_0^2+2ad,""d=\dfrac{v^2-v_0^2}{2a}""d=\dfrac{0-(30\ \dfrac{m}{s})^2}{2\times(-2.0\ \dfrac{m}{s^2})}=225\ m."